OpenStudy (anonymous):

Choose c so that the line y = x is tangent to the parabola y = x2 + C. They have the same slope where they touch.

OpenStudy (anonymous):

@Mohsin007 Any idea where to start?

OpenStudy (anonymous):

by derivatives i think

OpenStudy (anonymous):

Yes, from the key word "Tangent" :) Will you show your derivative here ?

OpenStudy (anonymous):

dy/dx=2x

OpenStudy (anonymous):

Now what's the relation between y' and the tangent line y = x?

OpenStudy (anonymous):

equal?

OpenStudy (anonymous):

Yep, but equal with ....?

OpenStudy (anonymous):

2x=1 x=1/2?

OpenStudy (anonymous):

ok ok i got it..:) at point of tangency x=x^2+c 1/2=1/4+c c=1/4

OpenStudy (anonymous):

1=2x

OpenStudy (anonymous):

why x=0?